if a polynomial P with integer coefficients has three distinct integer zeroes, then show that P(n) not equal to 1 for any integer n.

Grade:11

2 Answers

Aman Bansal
12 years ago

Dear Aritra,

The question is incorrect,

There are so many examples to prove it incorrect such as :

let P = (x-1)(x-2)(x-3)* (1/6 )

its zeroes are 1,2,3 which are integers and its value at 4(an integer) is 1.

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

 

Ashwin Muralidharan IIT Madras
12 years ago

Hi Aritra,

 

The above example P(x) = 1/6*(x-1)(x-2)(x-3) does not have integer co-efficients.

 

Let P(x) = (x-a)(x-b)(x-c)*Q(x). where A,a,b,c are integers with a,b,c being distinct. Q(x) is any polynomial with integer co-efficients.

Now for any integer n, consider P(n).

P(n) = (n-a)(n-b)(n-c)Q(n).

In the above expression Q(n) is any integer.

 

Assume P(n) = 1 for some n.

Since all are integers, only possibility is |n-a| = |n-b| = |n-c| = |Q(n)| = 1.

Consider only |n-a|=|n-b|=|n-c|=1 ------------------ (1)

Expression (1) can never be true for distinct integers a,b,c as explained below.

 

|n-a|=|n-b|

Only possibility n-a = b-n, ie a+b=2n.

similarly b+c=2n.

Subtraction a-c=0 or a=c --------- (contradiction).

 

Hence P(n) can never be equal to 1 for any integer n.

 

Regards,

Ashwin (IIT Madras).